# Preference-Driven Multi-Objective Combinatorial Optimization with Conditional Computation

### v2

@@ Appendix C MOCOP Solvers with POCCO
Existing nerual MOCOP solvers are based on a transformer-based architecture that consists of an encoder and a decoder. The encoder is used to generate embeddings for all nodes based on the instance $\mathcal{G}$ and the weight vector $\lambda$. The decoder is used to decode a sequence of actions based on these embeddings in an iterative fashion. To demonstrate the versatility of our POCCO, we integrate it into two SOTA neural methods, WE-CA [6] and CNH [14], yielding POCCO-W and POCCO-C.
@@ C.1 POCCO-W
Specifically, the conditional attention sublayer first condition node embeddings on the weight embedding via a feature-wise affine transform:
@@
where $W^{\gamma}$ and $W^{\beta}$ are trainable matrices; $\circ$ denotes element-wise multiplication. Then, the weight and node embeddings are updated via the Multi-Head Attention (MHA) mechanism with 8 heads and an Add & Norm, as follows:
@@
Given the eventual node embeddings $\{h_{i}\}_{i=0}^{n}$ and weight embedding $h_{\lambda}$ output by the encoder, the decoder autoregressively computes the probability of node selection over $T$ steps. At decoding step $t\in\{1,\dots,T\}$, the advanced context vector $h_{c}$ is produced by a MHA layer with 8 heads based on the context embedding $v_{c}$ and eventual node embeddings as follows:
@@
where $C$ is set to 50 and $W^{K}$ is a learnable weight matrix. Finally, the node-selection probability for the scalarized subproblem is given by:
@@ C.2 POCCO-C
Besides, POCCO-C employ the sinusoidal encoding based on sine and cosine functions to yield the PSE as follows,
@@
where $\xi(\xi=n+1)$ and $i$ ($i\in\left\{0,\cdots,63\right\}$) mean the problem size and dimension. The resulting d-dimensional PSEs are then processed using two linear layers with trainable matrices $W_{\xi 1}\in\mathbb{R}^{d\times 2d}$ and $W_{\xi 2}\in\mathbb{R}^{2d\times d}$. POCCO-C inject the size information by adding the results from PSE to $\{h_{i}\}_{i=0}^{n}$ from the encoder such that,
@@ C.3 Gating Mechanism
We employ subproblem-level gating, routing each context vector $h_{c}$ independently to a subset of experts. Let $d$ be the hidden dimension and $W_{G}\in\mathbb{R}^{d\times(m+1)}$ the trainable gating weights in POCCO. Given a batch of context vectors $X=\{h_{c}^{b}\}_{b=1}^{B},X\in\mathbb{R}^{B\times d}$, where $B$ is batch size, the gating network computes the score matrix $H\;=\;X\,W_{G}\;\in\;\mathbb{R}^{B\times(m+1)}$. Subproblem $b$ (with context vetor $h_{c}^{b}$) is routed to expert $E_{j}$ according to the $i$-th row of $H$. Each subproblem selects the top $k$ experts by score. For example, $k=2$ routing sends each subproblem to the two highest-scoring experts.
@@ Appendix D Training Algorithm
5:     $y_{j,p}^{b}\sim$ PairwisePreference$(1_{[\pi^{j,b}\lessdot\pi^{p,b}]}),\quad\forall j,p\in\{1,\cdots K\},\ \forall b\in\left\{1,\cdots,B\right\}$
@@
7:     $\theta\leftarrow$ ADAM($\theta,\nabla_{\theta}\mathcal{L}(\theta)$)
@@ Appendix F Hypervolume
Hypervolume (HV) is a widely-used indicator to evaluate approximate Pareto solutions to MOCOPs. Formally, the HV for a set of solutions $\mathcal{P}$ is defined as the volume of the subspace, which is weakly dominated by the solutions in $\mathcal{P}$ and bounded by a reference point $r^{*}$, such that,
@@
where $\zeta^{\kappa}$ denotes the Lebesgue measure on the $\kappa$-dimensional space, i.e., the volume for a $m$-dimensional subspace [55]. Since the range of objective values varies among different problems, we report the normalized HV $\bar{H}(\mathcal{P})=\text{HV}(\mathcal{P})/\prod_{i=1}^{\kappa}|r^{*}_{i}-z_{i}|$, where the ideal point $z=(z_{1},\dots,z_{\kappa})$ satisfies $z_{i}<\min\bigl{\{}f_{i}(\pi)\mid\pi\in\mathcal{P}\bigr{\}}\bigl{(}\text{or }z_{i}>\max\{f_{i}(\pi)\mid\pi\in\mathcal{P}\}\text{ for maximization}\bigr{)},\forall\,i\in\{1,\dots,\kappa\}$. All methods share the same $r^{*}$ and $z$ for a MOCOP, as given in Table 3, and we report the average $\bar{H}(\mathcal{P})$ over all test instances in this paper.
@@ Appendix G Instance Augmentation
To further improve the performance of POCCO at the inference stage, we apply the instance augmentation proposed in [30], which is also used in PMOCO [39]. The rationale of instance augmentation is that an instance of Euclidean VRPs can be transformed into different ones, that share the same optimal solution, e.g., by flipping coordinates for all nodes in an instance. Given a coordinate $(x,y)$ in a VRP, there are eight simple transformations, i.e., $(x^{\prime},y^{\prime})$ = $(x,y);(y,x);(x,1-y);(y,1-x);(1-x,y);(1-y,x)$; $(1-x,1-y);(1-y,1-x)$. In our paper, we adopt these transformations for each objective, respectively. Hence, we could have 8 transformations for Bi-CVRP (since there is only one coordinate for each node), $648^{2}\!=\!$ 64 transformations for Bi-TSP and $5128^{3}\!=\!$ 512 transformations for Tri-TSP, respectively.
Appendix H Detailed Results on Benchmark Instances
@@
Appendix I Experimental Results on the Larger Probblem Sizes
Table 5: Performance on Bi-TSP150 and Bi-TSP200 Instances
Bi-TSP150 Bi-TSP200
@@ Appendix I Experimental Results on the Larger Probblem Sizes
Appendix J Hyperparameter Study
Effects of the $\beta$. Fig.5 shows how the temperature parameter $\beta$ in the preference learning loss affects performance (HV) across four benchmark tasks. A moderate value of $\beta$ consistently yields the best results. For the three bi-objective problems (Bi-TSP100, MOCVRP100, Bi-KP100) performance peaks around $\beta=3.5$; larger or smaller values provide no additional benefit, and in Bi-KP100 an overly large $\beta=5$ sharply degrades HV. For the tri-objective problem (Tri-TSP100), performance improves up to $\beta=4.5$ and then declines slightly. Overall, these trends justify the default settings adopted in our experiments: $\beta=3.5$ for bi-objective tasks and $\beta=4.5$ for tri-objective tasks.
@@
Effects of the number of CCO layers. The CCO block sits in the decoder, whose sequence length T grows with problem size. Each additional CCO layer therefore adds a full set of gating and expert computations at every decoding step. Hence, stacking too many CCO layers can inflate runtime disproportionately. Table 6 reports the trade-off. Using two CCO layers yields the highest HV on both Bi-TSP100 and Bi-TSP200, but increases inference time by roughly 60% and 130 %, respectively. A third layer further slows inference while slightly reducing HV. We therefore adopt a single-layer CCO as the default, which preserves most of the performance gain while keeping computation modest. For larger instances or latency-sensitive applications, the one-layer setting offers a favorable balance between solution quality and speed.
@@
Effects of $\operatorname{Top}k$. Fig.6 plots validation HV on Bi-TSP100 for $k\!\in\!\{1,2,3\}$. Selecting two experts per token ($k$=2) converges fastest and attains the highest final HV. A single expert ($k$=1) limits model capacity, leading to slower early progress and a slightly lower plateau. Choosing three experts ($k$=3) increases computation relative to $k$=2 yet offers no benefit and even marginally reduces HV, likely because the additional expert dilutes specialization and weakens sparsity. We therefore adopt $k$=2 as the default, balancing performance and efficiency.
@@
Appendix K Experimental Results of Scalarized Subproblems
@@
Appendix L Summary of Decomposition-based Neural MOCOP Solvers
@@
Appendix M Broader Impacts
@@
Appendix N Licenses for Existing Assets
The used assets in this work are listed in Table 9, which are all open-source for academic research. We will release our source code with the MIT License.
@@
DRL-MOA [34] MIT License Evaluation
MDRL [68] MIT License Evaluation
EMNH [7] MIT License Evaluation

### v4

@@ Appendix C MOCOP Solvers with POCCO
Existing neural MOCOP solvers are based on a transformer-based architecture that consists of an encoder and a decoder. The encoder is used to generate embeddings for all nodes based on the instance $\mathcal{G}$ and the weight vector $\lambda$. The decoder is used to decode a sequence of actions based on these embeddings in an iterative fashion. To demonstrate the versatility of our POCCO, we integrate it into two SOTA neural methods, WE-CA [6] and CNH [14], yielding POCCO-W and POCCO-C.
@@ C.1 POCCO-W
Specifically, the conditional attention sublayer first conditions node embeddings on the weight embedding via a feature‐wise affine transform:
@@
where $W^{\gamma}$ and $W^{\beta}$ are trainable matrices; $\circ$ denotes element‐wise multiplication. Then, the weight and node embeddings are updated via the Multi‐Head Attention (MHA) mechanism with 8 heads and an Add & Norm, as follows:
@@
Given the eventual node embeddings $\{h_{i}\}_{i=0}^{n}$ and weight embedding $h_{\lambda}$ output by the encoder, the decoder autoregressively computes the probability of node selection over $T$ steps. At decoding step $t\in\{1,\dots,T\}$, the advanced context vector $h_{c}$ is produced by an MHA layer with 8 heads based on the context embedding $v_{c}$ and eventual node embeddings as follows:
@@
where $C$ is set to 50 and $W^{K}$ is a learnable weight matrix. Finally, the node‐selection probability for the scalarized subproblem is given by:
@@ C.2 POCCO-C
Besides, POCCO-C employs the sinusoidal encoding based on sine and cosine functions to yield the PSE as follows,
@@
where $\xi(\xi=n+1)$ and $i$ ($i\in\left\{0,\cdots,63\right\}$) mean the problem size and dimension. The resulting d-dimensional PSEs are then processed using two linear layers with trainable matrices $W_{\xi 1}\in\mathbb{R}^{d\times 2d}$ and $W_{\xi 2}\in\mathbb{R}^{2d\times d}$. POCCO-C injects the size information by adding the results from PSE to $\{h_{i}\}_{i=0}^{n}$ from the encoder, such that,
@@ C.3 Gating Mechanism
Our CCO block employs a subproblem-level gating mechanism to dynamically route each context vector $h_{c}$ to a subset of experts from a pool consisting of 4 FF experts and 1 ID expert (thus $m=4$). Let $d$ be the hidden dimension and $W_{G}\in\mathbb{R}^{d\times(m+1)}$ denote the trainable gating weight matrix. Given a batch of context vectors $X=\{h_{c}^{b}\}_{b=1}^{B}\in\mathbb{R}^{B\times d}$, where $B$ is the batch size, the gating network computes a score matrix: $H=X\cdot W_{G}\in\mathbb{R}^{B\times(m+1)}$, where each element $H_{b}^{j}$ represents the affinity or preference (score) of subproblem $b$ towards expert $j$.
For each subproblem, the router selects the Top-$k$ highest-scoring experts from the score vector $H_{b}^{j}$. These selected experts are then activated, and their outputs are weighted by a softmax-normalized version of their scores. In our POCCO with $k=2$, each subproblem is routed to its two highest-scoring experts.
This Top-$k$ routing plays a critical role by reducing computation, as only $k$ experts are activated per subproblem (i.e., sparse activation widely used in MoE structure [69]). It also encourages expert specialization, since different subproblems (defined by context and weight vectors) tend to activate different subsets of experts. Additionally, it allows the inclusion of a parameter-free ID expert, which is often selected when minimal transformation is needed. This enables the router to learn when it is appropriate to leave a representation unchanged.
@@ Appendix D Training Algorithm
4:  $\pi^{j,b}\!\simSampleSolutions$($p_{\theta}(\cdot|\mathcal{G}_{b},\lambda_{b})$), $\forall j\in\left\{1,\cdots,K\right\},\quad\forall b\in\left\{1,\cdots,B\right\}$
@@
[…]
@@ Appendix F Hypervolume
Hypervolume (HV) is a widely used indicator to evaluate approximate Pareto solutions to MOCOPs. Formally, the HV for a set of solutions $\mathcal{P}$ is defined as the volume of the subspace, which is weakly dominated by the solutions in $\mathcal{P}$ and bounded by a reference point $r^{*}$, such that,
@@
where $\zeta^{\kappa}$ denotes the Lebesgue measure on the $\kappa$-dimensional space, i.e., the volume for a $\kappa$-dimensional subspace [54]. Since the range of objective values varies among different problems, we report the normalized HV $\bar{H}(\mathcal{P})=\text{HV}(\mathcal{P})/\prod_{i=1}^{\kappa}|r^{*}_{i}-z_{i}|$, where the ideal point $z=(z_{1},\dots,z_{\kappa})$ satisfies $z_{i}<\min\bigl\{f_{i}(\pi)\mid\pi\in\mathcal{P}\bigr\}\bigl(\text{or }z_{i}>\max\{f_{i}(\pi)\mid\pi\in\mathcal{P}\}\text{ for maximization}\bigr),\forall\,i\in\{1,\dots,\kappa\}$. All methods share the same $r^{*}$ and $z$ for an MOCOP, as given in Table 4, and we report the average $\bar{H}(\mathcal{P})$ over all test instances in this paper.
@@ Appendix G Instance Augmentation
To further improve the performance of POCCO at the inference stage, we apply the instance augmentation proposed in [30], which is also used in PMOCO [38]. The rationale of instance augmentation is that an instance of Euclidean VRPs can be transformed into different ones that share the same optimal solution, e.g., by flipping coordinates for all nodes in an instance. Given a coordinate $(x,y)$ in a VRP, there are eight simple transformations, i.e., $(x^{\prime},y^{\prime})$ = (x,y);(y,x);(x,1y);(y,1x);(1x,y);(1y,x)(x,y);(y,x);(x,1-y);(y,1-x);(1-x,y);(1-y,x); (1x,1y);(1y,1x)(1-x,1-y);(1-y,1-x). In our paper, we adopt these transformations for each objective, respectively. Hence, we could have 8 transformations for Bi-CVRP (since there is only one coordinate for each node), $8^{2}\!=\!64$ transformations for Bi-TSP, and $8^{3}\!=\!512$ transformations for Tri-TSP, respectively.
Appendix H Impact of Neutralizing Non-Dominated Pairs
Our method adopts a decomposition-based framework, where the MOCOP is divided into a set of scalarized subproblems, each associated with a weight vector. During training, the goal is to find high-quality solutions with respect to each subproblem, rather than globally exploring the entire Pareto front at once. In this context, preference learning operates at the subproblem level. We acknowledge, however, that sampled solutions may occasionally be mutually non-dominated in the original objective space. To investigate the effect of this scenario, we conducted an additional experiment where preference signals for such non-dominated pairs were set to 0, effectively treating them as equivalent. The results, shown in Table 5, indicate that this modification (NDS) significantly degrades model performance across all Bi-TSP instances.
This observation suggests that introducing indifference signals for non-dominated pairs hinders the learning process. One reason is that it reduces the amount of effective preference supervision per subproblem, especially in early training stages when many solutions are of similar quality. In contrast, using scalarized values provides a consistent and differentiable training signal aligned with the decomposition objective. Therefore, although scalarized comparisons may not reflect global Pareto dominance in every case, they remain well-justified and necessary within the decomposition-based learning paradigm, and do not appear to bias the overall Pareto front approximation based on our empirical findings.
Table 5: Performance on Bi-TSP instances when neutralizing non-dominated pairs.
Problem Size POCCO-W HV NDS HV
Bi-TSP20 0.6275 0.5516
Bi-TSP50 0.6411 0.5160
Bi-TSP100 0.7055 0.5601
Bi-TSP150 0.7033 0.5494
Bi-TSP200 0.7371 0.5809
Appendix I Detailed Results on Benchmark Instances
@@
Table 7: Performance on Bi‐TSP150 and Bi‐TSP200 Instances
Bi‐TSP150 Bi‐TSP200
@@
Table 8: HV on large Bi-TSP instances.
Problem Size WE-CA POCCO-W
Bi-TSP300 0.7441 0.7458
Bi-TSP500 0.7476 0.7592
Bi-TSP1000 0.7186 0.7400
Table 9: HV on large MOKP instances.
Problem Size WE-CA POCCO-W
Bi-KP300 0.6000 0.5947
Bi-KP500 0.4244 0.5658
Bi-KP1000 0.2439 0.8408
Appendix J Experimental Results on the Larger Problem Sizes
Furthermore, to demonstrate scalability, we include experimental results on even larger problems, including Bi-TSP with 300/500/1000 nodes and MOKP with 300/500/1000 items. These results are summarized in Table 8 and Table 9. As shown, POCCO-W consistently outperforms the baseline WE-CA in most large-scale cases.
Appendix K Comparison with DPO
We have added a comparative experiment between our preference learning (PL) method and the recent DPO objective, as shown in Table 10. Since DPO requires two models, i.e., a policy model and a reference model, we use the same architecture and initialization for both, with the reference model updated from the previous epoch. This setup significantly increases training time for DPO (122h vs. 36h for POCCO-W).
As shown in Table 10, while DPO achieves slightly better performance on the smallest instance (MOTSP20), it is consistently outperformed by POCCO-W on larger instances (MOTSP50–200). These results suggest that our PL method offers a more favorable trade-off in terms of both training efficiency and solution quality, especially for large-scale problems.
Table 10: HV of different PL methods on Bi-TSP instances.
Problem Size POCCO-W DPO
Bi-TSP20 0.6275 0.6276
Bi-TSP50 0.6411 0.6400
Bi-TSP100 0.7055 0.7030
Bi-TSP150 0.7033 0.6998
Bi-TSP200 0.7371 0.7331
Appendix L Hyperparameter Study
Effects of the $\beta$. Fig.5 shows how the temperature parameter $\beta$ in the preference learning loss affects performance (HV) across four benchmark tasks. A moderate value of $\beta$ consistently yields the best results. For the three bi-objective problems (Bi-TSP100, MOCVRP100, Bi-KP100), performance peaks around $\beta=3.5$; larger or smaller values provide no additional benefit, and in Bi-KP100, an overly large $\beta=5$ sharply degrades HV. For the tri-objective problem (Tri-TSP100), performance improves up to $\beta=4.5$ and then declines slightly. Overall, these trends justify the default settings adopted in our experiments: $\beta=3.5$ for bi-objective tasks and $\beta=4.5$ for tri-objective tasks.
@@
Effects of the number of CCO layers. The CCO block sits in the decoder, whose sequence length T grows with problem size. Each additional CCO layer, therefore, adds a full set of gating and expert computations at every decoding step. Hence, stacking too many CCO layers can inflate the runtime disproportionately. Table 11 reports the trade-off. Using two CCO layers yields the highest HV on both Bi-TSP100 and Bi-TSP200, but increases inference time by roughly 60% and 130 %, respectively. A third layer further slows inference while slightly reducing HV. We therefore adopt a single-layer CCO as the default, which preserves most of the performance gain while keeping computation modest. For larger instances or latency-sensitive applications, the one-layer setting offers a favorable balance between solution quality and speed.
@@
Effects of the number of experts. We have added an ablation study to analyze the impact of the number of experts in Table 12. Specifically, POCCO-W (5E) refers to the original model presented in the paper, which includes 4 feedforward (FF) experts and 1 parameter-free identity (ID) expert. We additionally evaluate four variants:
  1. 1.
    POCCO-W (3E): 2 FF + 1 ID;
  2. 2.
    POCCO-W (9E): 8 FF + 1 ID;
  3. 3.
    POCCO-W (9E_2D): 8 FF + 1 ID, trained on twice the amount of data;
  4. 4.
    POCCO-W (17E): 16 FF + 1 ID.
All variants are trained under the same settings as POCCO-W (5E) for a fair comparison, except for POCCO-W (9E_2D), which is trained on more data.
As shown in Table 12, all POCCO-W variants outperform the backbone model WE-CA, confirming the effectiveness of the expert-based architecture. Moreover, POCCO-W (3E) and POCCO-W (5E) achieve better overall performance than POCCO-W (9E) and POCCO-W (17E), the latter of which requires additional data scaling to realize performance gains. To strike a better balance between computational cost and solution quality, we select POCCO-W (5E) as our default model.
Table 12: Effects of the number of experts.
Problem Size WE-CA POCCO-W (3E) POCCO-W (5E) Exp_num (9E) Exp_num (9E_2D) Exp_num (17E)
Bi-TSP20 0.6270 0.6275 0.6275 0.6275 0.6275 0.6275
Bi-TSP50 0.6392 0.6410 0.6411 0.6408 0.6411 0.6408
Bi-TSP100 0.7034 0.7057 0.7055 0.7050 0.7053 0.7050
Bi-TSP150 0.7008 0.7035 0.7033 0.7026 0.7031 0.7027
Bo-TSP200 0.7346 0.7363 0.7371 0.7364 0.7370 0.7361
Effects of $\operatorname{Top}k$. Fig.6 plots validation HV on Bi-TSP100 for $k\!\in\!\{1,2,3\}$. Selecting two experts per token (kk=2) converges fastest and attains the highest final HV. A single expert (kk=1) limits model capacity, leading to slower early progress and a slightly lower plateau. Choosing three experts (kk=3) increases computation relative to kk=2 yet offers no benefit and even marginally reduces HV, likely because the additional expert dilutes specialization and weakens sparsity. We therefore adopt kk=2 as the default, balancing performance and efficiency.
@@
Appendix M Experimental Results of Scalarized Subproblems
@@
Appendix N Comparison with Tchebycheff Scalarization
Intuitively, our POCCO framework is applicable to any MOCOP where the objectives can be scalarized using decomposition-based techniques, including weighted-sum (WS), Tchebycheff (TCH), and penalty-based boundary intersection (PBI). This is because the preference signal between two solutions is determined solely by the ordering of their scalarized objective values, regardless of the specific scalarization method used.
To empirically demonstrate this generality, we added an experiment comparing POCCO-W trained with WS and TCH scalarization methods. As shown in Table 14, WS consistently outperforms TCH on Bi-TSP50 to Bi-TSP200, indicating that WS offers more robust performance across different problem sizes. This observation is also consistent with findings in [6], which report that WS—despite its simplicity—often outperforms TCH in the studied MOCOP setting. In general, however, WE-CA (TCH) performs worse than WE-CA (WS), which achieves a score of 0.6392, and is further outperformed by POCCO-W (TCH), which achieves 0.6395 on Bi-TSP50.
Table 14: HV of Different Decomposition Approaches on Bi-TSP Instances.
Problem Size WE-CA (WS) POCCO-W (WS) POCCO-W (TCH)
Bi-TSP50 0.6392 0.6411 0.6395
Bi-TSP100 0.7034 0.7055 0.7031
Bi-TSP150 0.7008 0.7033 0.6998
Bi-TSP200 0.7346 0.7371 0.7329
Appendix O GPU Memory Usage
Table 15: Comparison of GPU Memory Usage.
Problem Size WE-CA CNH POCCO-W
Bi-TSP50 1051 MB 1077 MB 1849 MB
MOCVRP50 1790 MB 1501 MB 4522 MB
Bi-KP50 916 MB 916 MB 2417 MB
Tri-TSP50 1052 MB 1077 MB 2383 MB
We present a comparison of GPU memory usage between POCCO-W (ours), the backbone model (WE-CA), and CNH in Table 15. All models are trained across problem instances with size $n=50$. As shown in Table 15, POCCO-W incurs higher GPU memory usage, approximately doubling that of the baselines WE-CA and CNH. However, this increased resource cost is accompanied by significant improvements in solution quality, as demonstrated in our experimental results.
Appendix P Summary of Decomposition-based Neural MOCOP Solvers
@@
Appendix Q Broader Impacts
@@
Appendix R Licenses for Existing Assets
The used assets in this work are listed in Table 17. Where applicable, we reference publicly available implementations for evaluation or reproduction purposes. Our source code is released under the MIT License.
@@
DRL-MOA [34] No license (assumed all rights reserved) Evaluation
MDRL [67] No license (assumed all rights reserved) Evaluation
EMNH [7] No license (assumed all rights reserved) Evaluation
PMOCO [38] MIT License Evaluation